has to be answered in full equation form..
alex is twice as old as chris. twelve years ago alex was three times as old as chris. how old are they now?

Answer by swincher4391(1107)

a = 2c
a-12 = 3(c-12)
2c-12 = 3c -36
-c = -24
c = 24
Chris is 24.
a = 2(24)
Alex is 48.